# Dual-component stellar assembly histories in local elliptical galaxies via MUSE

### v1

@@ 5.3 Star formation histories: stellar assembly across time
Even from a cursory look, it would appear that the ellipticals in our sample fit the description of “red and dead” retired galaxies commonly referenced in literature. This impression is supported by the cumulative star formation histories (Fig. 7, 8, and 9), which reveal a significant early and rapid assembly of stellar mass for the inner component, heavily influencing its mass and luminosity. Similarly, the outer component exhibits a concentrated period of star formation in its early life, resulting in the majority of its stellar mass and luminosity. We surmise that the inner and outer components represent the in-situ and ex-situ components respectively, motivated by similar works described in Huang et al. (2013b, a); Spavone et al. (2017, 2021). Following the initial burst of star formation, the mass assembly histories of both components allude to rather quiescent ellipticals where there has been no significant build-up of stellar mass over roughly the last 8 Gyr. The star formation histories that account for the luminosity contributions by the stellar populations in each component depict a slightly nuanced picture. The majority of the light does indeed originate from the initial rapid star formation episode shortly after the Big Bang, and we find no indications of younger stellar populations contributing to the stellar mass or luminosity of any of the galaxies.
@@
The star formation histories of the outer components are relatively more extended than those of the inner components, with the peak of star formation and mass assembly shifted towards relatively recent epochs. However, from Table 8, for galaxies J020536 and J205050, the contribution of stars younger than 3 Gyr to the stellar mass or luminosity is still entirely negligible. The fraction of intermediate-age stars is however quite substantial next to the oldest stars, suggesting this population either formed in-situ later in time or was accreted as an ex-situ population through dry mergers with relatively younger satellite galaxies aged between 4 and 8 Gyr old. Such a blend of old and intermediate-age stars were measured for dwarf elliptical galaxies in the Virgo and Fornax galaxy clusters (Geha et al., 2003; Michielsen et al., 2008), and in the Local Group (Geha et al., 2015). The unperturbed nature of the quiescence of these elliptical galaxies provide additional evidence that they are classical “red and dead” objects in the local Universe, where both in-situ and ex-situ star formation have ceased several Gyr ago. In that vein, the following segments of the discussion outline other studies in literature that have investigated the assembly histories of elliptical galaxies through various observations and techniques.
@@
Lacerna et al. (2020) analysed the integrated star formation histories of 251 classical elliptical galaxies in MaNGA. They find that these galaxies had built up 70% of their stellar masses between 8 and 11 Gyr ago, and 90% of their masses by 5 Gyr ago at the latest. In conjunction with their results on the stellar population gradients (we refer the reader to Sect. 5.2), they predict that in the earliest stages of elliptical galaxy formation, star formation and the resulting stellar populations were relatively uniform across both the inner and outer regions. It was only in subsequent epochs that the outer regions started forming their stars after the inner regions did. Moreover, by analysing the specific star formation rates of the galaxies, they conclude that, following an initial phase of substantial homogeneous star formation, classical ellipticals continue to grow from the inside out and are subsequently quenched in the same manner. This was naturally explained on the basis of an intense burst of significant star formation, followed by rapid gas consumption, leading to a swift quenching of star formation. Nevertheless, they recognise the role of dry mergers where ex-situ stars can be accreted on to the main elliptical, and continue its stellar mass assembly. Despite the methodology and sample used being entirely different to theirs, the inferences we present are remarkably similar in the context of the two-phase formation scenario for early-type galaxies. However, it is important to note that the fossil-record method does not directly characterise a given stellar population as having formed via in-situ or ex-situ processes.
@@
A study by González Delgado et al. (2017) employs the fossil-record method to reconstruct the spatially resolved 2D SFHs of galaxies spanning a wide variety of morphologies in the CALIFA survey. Similar to Table 8 in this work, they report the fractions of the stars in different stellar age bins which contribute to the luminosity of the galaxy. They find that % of stars younger than 1 Gyr contribute to the light from the outer regions of ellipticals, while the older stars with ages over 4 Gyr dominate with % of the populations. In the inner regions, that fraction increases to 96% for the older stars, and 1% for the younger population. We note that this is slightly in contrast to our sample, which reveals no traces of young stellar populations in the inner components. The contributions to the stellar mass follow the same pattern in the inner and outer regions. Furthermore, they find evidence of an extended phase of star formation since 4 Gyr ago, and signs of the outer regions of ellipticals that are actively forming stars after the quiescent phase.
@@
Another sample of nearby ETGs from the survey (Cappellari et al., 2011) were investigated in McDermid et al. (2015), where they measure the spatially resolved stellar populations and integrated star formation histories within of the galaxies. Assessing this region within across the galaxies in our sample (see Table 7), we note that this would result in a loss of information about the outskirts or the outer component, which extend to at least twice the from the single Sérsic model. They find that roughly half the present-day stellar mass was already assembled as early as 2 Gyr after the Big Bang. The most massive present-day ETGs with host these old stars, and had built up 90% of the stellar mass by , which corresponds to a lookback time of approximately 10.5 Gyr ago under the cosmology defined in Sect. 1. They attribute their results to the picture where the old and metal-rich ETGs were formed from an early dissipative process at high redshifts, which is followed by a phase of quenching. Subsequently, stellar accretion takes over as the dominant mechanism of stellar mass growth, aligning with the two-phase scenario of formation.
@@
The importance of ex-situ stars that have been accreted by early-type galaxies in their later stages has been demonstrated by observations and simulations, particularly with the IFS surveys. The metallicity and age distributions and profiles are often exploited to estimate the fraction of ex-situ accreted stars in an ETG. For instance, Oyarzún et al. (2019) measure the ex-situ stellar mass fractions of ETGs in the MaNGA survey. They find that for ETGs of , ex-situ stars make up to of their stellar mass, at a galactocentric distance of . Similarly, using a combination of full-spectral fitting and galaxy evolution models, Davison et al. (2021) identify ex-situ stars in ETGs observed with MUSE, and report that the fraction of accreted stars increases with galactocentric radius, with values ranging between and 100% at . Spavone et al. (2021) present deep imaging of three massive galaxies from the VEGAS survey, combined with MUSE observations. This study correlates the structural components from imaging with the kinematics and stellar populations from IFS analysis, to explore the mass assembly histories of these galaxies. They find ex-situ mass fractions of 77%, 86%, and 89%, indicating that accretion is the primary contributor to the stellar mass. Our results align with the ranges reported by some of these studies, indicating that 50% of the stellar mass was accreted from external galaxies, forming a second “co-dominant” component in the context of a dual-component picture of elliptical galaxies.
@@ 6 Summary and conclusions
In summary, this study was primarily focused on the identification of the different components present in elliptical galaxies of intermediate stellar mass in the local Universe. For a sample of three galaxies observed with MUSE, we successfully modelled the surface brightness profiles adopting two Sérsic components, representing an inner and an outer component, using the spectro-photometric decomposition code BUDDI. After this, the full-spectral fitting software pPXF was used to fit the cleanly extracted spectra and estimate the mass-weighted and light-weighted stellar populations of each component, and subsequently their star formation histories. Spatially resolved 2D stellar population maps were constructed after binning the datacubes to achieve a sufficient minimum S/N threshold using the Voronoi tessellation method. We investigated the physical properties of each component to constrain the formation scenarios and the mode of stellar mass growth in local elliptical galaxies. While similar studies with photometric decomposition on data from imaging surveys have been conducted, this approach of using IFS to model multiple components in datacubes is relatively new and offers a different perspective of exploring elliptical galaxy formation and evolution. The main conclusions of our study are listed below:
@@
The best fit for the surface brightness profiles of the elliptical galaxies in this sample was achieved using two Sérsic components with distinct properties, referred to as the ”inner” and ”outer” components.
@@
The stellar population analysis of the inner and outer components and the spatial maps reveal an inside-out formation and growth mode in ellipticals. The majority of the stellar populations are older than 8 Gyr for both components (100% for the inner component, and for the outer component), from both the mass and light-weighted ages. The inner components are generally more metal-enhanced compared to the outer components.
@@
The reconstructed SFHs highlight an inner component that has formed and assembled the bulk of its stars in the earliest epochs of the Universe, before . This component is expected to have formed from interlinked physical processes: rapid dissipative collapse and major mergers, forming a central core hosting the oldest stellar populations. The outer component, however, aligns more with accretion events from past dry mergers, where old metal-poor stars have been deposited on the outskirts. This process continues to build the stellar mass of the ellipticals without having to undergo a new phase of star formation. The fraction of mass and luminosity contributed by the different stellar populations in the two components further confirms the two-phase assembly mode.
@@
The mass weights from pPXF additionally provide information about the formation ( and quenching ( lookback times, which show that half of the present-day stellar mass in our sample had already assembled Gyr ago, and had undergone quenching around Gyr ago at the latest. The light-weighted properties show similar trends, except , which indicates a more extended star formation history in the outer component, with 90% of the light contribution coming from lookback times as recent as Gyr ago.
@@
In conclusion, our method of decomposing the multiple components of elliptical galaxies has helped trace back the signatures of the physical processes involved in their formation and evolution across cosmic time. This serves as an initial pilot study, which can be extended to a more statistical sample such as the ellipticals in BUDDI-MaNGA (e.g. Jegatheesan et al., 2024) to better constrain the formation pathways. Possible future follow-up work involves capturing the extremely low-surface brightness stellar halos of elliptical galaxies and modelling their components (Johnston et al., 2018). Additionally, studying a more representative sample of local ellipticals, encompassing a range of masses from dwarf to giant ellipticals, variations in star formation activity, and in different environments, would provide a comprehensive understanding of the formation and evolution of these galaxies.

### v2

@@ 5.3 Star formation histories: Stellar assembly across time
Even from a cursory look, it would appear that the ellipticals in our sample fit the description of red and dead retired galaxies commonly referenced in the literature. This impression is supported by the cumulative SFHs (Fig. 7, 8, and 9), which reveal a significant early and rapid assembly of stellar mass for the inner component, heavily influencing its mass and luminosity. Similarly, the outer component exhibits a concentrated period of star formation in its early life, resulting in the majority of its stellar mass and luminosity. We surmise that the inner and outer components represent the in situ and ex situ components, respectively, motivated by similar works described in Huang et al. (2013b, a); Spavone et al. (2017, 2021). Following the initial burst of star formation, the mass assembly histories of both components allude to rather quiescent ellipticals where there has been no significant build-up of stellar mass over roughly the last 8 Gyr. The SFHs that account for the luminosity contributions by the stellar populations in each component depict a slightly nuanced picture. The majority of the light does indeed originate from the initial rapid star formation episode shortly after the Big Bang, and we find no indications of younger stellar populations contributing to the stellar mass or luminosity of any of the galaxies.
@@
The SFHs of the outer components are more extended than those of the inner components, with the peak of star formation and mass assembly shifted towards relatively recent epochs. However, from Table 8, for galaxies J020536 and J205050, the contribution of stars younger than 3 Gyr to the stellar mass or luminosity is still entirely negligible. The fraction of intermediate-age stars is however quite substantial next to the oldest stars, suggesting this population either formed in situ later in time or was accreted as an ex situ population through dry mergers with younger satellite galaxies aged between 4 and 8 Gyr old. Such a blend of old and intermediate-age stars were measured for dwarf elliptical galaxies in the Virgo and Fornax galaxy clusters (Geha et al., 2003; Michielsen et al., 2008), and in the Local Group (Geha et al., 2015). The unperturbed nature of the quiescence of these elliptical galaxies provide additional evidence that they are classical red and dead objects in the local Universe, where both in situ and ex situ star formation have ceased several gigayears ago. In that vein, the following segments of the discussion outline other studies in the literature that have investigated the assembly histories of elliptical galaxies through various observations and techniques.
@@
Lacerna et al. (2020) analysed the integrated SFHs of 251 classical elliptical galaxies in MaNGA. They find that these galaxies had built up 70% of their stellar masses between 8 and 11 Gyr ago, and 90% of their masses by 5 Gyr ago at the latest. In conjunction with their results on the stellar population gradients (we refer the reader to Sect. 5.2), they predict that in the earliest stages of elliptical galaxy formation, star formation and the resulting stellar populations were relatively uniform across both the inner and outer regions. It was only in subsequent epochs that the outer regions started forming their stars after the inner regions did. Moreover, by analysing the specific star formation rates of the galaxies, they conclude that, following an initial phase of substantial homogeneous star formation, classical ellipticals continue to grow from the inside out and are subsequently quenched in the same manner. This was naturally explained on the basis of an intense burst of significant star formation, followed by rapid gas consumption, leading to a swift quenching of star formation. Nevertheless, they recognise the role of dry mergers where ex situ stars can be accreted on to the main elliptical, and continue its stellar mass assembly. Despite the methodology and sample used being entirely different to theirs, the inferences we present are remarkably similar in the context of the two-phase formation scenario for ETGs. However, it is important to note that the fossil-record method does not directly characterise a given stellar population as having formed via in situ or ex situ processes.
@@
A study by González Delgado et al. (2017) employed the fossil-record method to reconstruct the spatially resolved 2D SFHs of galaxies spanning a wide variety of morphologies in the CALIFA survey. Similar to Table 8 in this work, they report the fractions of the stars in different stellar age bins that contribute to the luminosity of the galaxy. They find that % of stars younger than 1 Gyr contribute to the light from the outer regions of ellipticals, while the older stars with ages over 4 Gyr dominate with % of the populations. In the inner regions, that fraction increases to 96% for the older stars, and 1% for the younger population. We note that this is slightly in contrast to our sample, which reveals no traces of young stellar populations in the inner components. The contributions to the stellar mass follow the same pattern in the inner and outer regions. Furthermore, they find evidence of an extended phase of star formation since 4 Gyr ago, and signs of the outer regions of ellipticals that are actively forming stars after the quiescent phase.
@@
Another sample of nearby ETGs from the survey (Cappellari et al., 2011) were investigated in McDermid et al. (2015), where they measure the spatially resolved stellar populations and integrated SFHs within of the galaxies. Assessing this region within across the galaxies in our sample (see Table 7), we note that this would result in a loss of information about the outskirts or the outer component, which extend to at least twice the from the single Sérsic model. They find that roughly half the present-day stellar mass was already assembled as early as 2 Gyr after the Big Bang. The most massive present-day ETGs with host these old stars, and had built up 90% of the stellar mass by , which corresponds to a look-back time of approximately 10.5 Gyr ago under the cosmology defined in Sect. 1. They attribute their results to the picture where the old and metal-rich ETGs were formed from an early dissipative process at high redshifts, which is followed by a phase of quenching. Subsequently, stellar accretion takes over as the dominant mechanism of stellar mass growth, aligning with the two-phase scenario of formation.
@@
The importance of ex situ stars that have been accreted by ETGs in their later stages has been demonstrated by observations and simulations, particularly with the IFS surveys. The metallicity and age distributions and profiles are often exploited to estimate the fraction of ex situ accreted stars in an ETG. For instance, Oyarzún et al. (2019) measure the ex situ stellar mass fractions of ETGs in the MaNGA survey. They find that for ETGs of , ex situ stars make up to of their stellar mass, at a galactocentric distance of . Similarly, using a combination of full-spectral fitting and galaxy evolution models, Davison et al. (2021) identify ex situ stars in ETGs observed with MUSE, and report that the fraction of accreted stars increases with galactocentric radius, with values ranging between and 100% at . Spavone et al. (2021) present deep imaging of three massive galaxies from the VEGAS survey, combined with MUSE observations. This study correlates the structural components from imaging with the kinematics and stellar populations from IFS analysis, to explore the mass assembly histories of these galaxies. They find ex situ mass fractions of 77%, 86%, and 89%, indicating that accretion is the primary contributor to the stellar mass. Our results align with the ranges reported by some of these studies, indicating that 50% of the stellar mass was accreted from external galaxies, forming a second co-dominant component in the context of a dual-component picture of elliptical galaxies.
@@ 6 Summary and conclusions
In summary, this study was primarily focused on the identification of the different components present in elliptical galaxies of intermediate stellar mass in the local Universe. For a sample of three galaxies observed with MUSE, we successfully modelled the surface brightness profiles adopting two Sérsic components, representing an inner and an outer component, using the spectro-photometric decomposition code BUDDI. After this, the full-spectral fitting software pPXF was used to fit the cleanly extracted spectra and estimate the mass-weighted and light-weighted stellar populations of each component, and subsequently their SFHs. Spatially resolved 2D stellar population maps were constructed after binning the datacubes to achieve a sufficient minimum S/N threshold using the Voronoi tessellation method. We investigated the physical properties of each component to constrain the formation scenarios and the mode of stellar mass growth in local elliptical galaxies. While similar studies with photometric decomposition on data from imaging surveys have been conducted, this approach of using IFS to model multiple components in datacubes is relatively new and offers a different perspective of exploring elliptical galaxy formation and evolution. The main conclusions of our study are listed below:
@@
The best fit for the surface brightness profiles of the elliptical galaxies in this sample was achieved using two Sérsic components with distinct properties, referred to as the inner and outer components.
@@
The stellar population analysis of the inner and outer components and the spatial maps reveal an inside-out formation and growth mode in ellipticals. The majority of the stellar populations are older than 8 Gyr for both components (100% for the inner component, and for the outer component), from both the mass and light-weighted ages. The inner components are generally more metal-enhanced than the outer ones.
@@
The reconstructed SFHs highlight an inner component that has formed and assembled the bulk of its stars in the earliest epochs of the Universe, before . This component is expected to have formed from interlinked physical processes: rapid dissipative collapse and major mergers, forming a central core hosting the oldest stellar populations. The outer component, however, aligns more with accretion events from past dry mergers, in which old metal-poor stars have been deposited on the outskirts. This process continues to build the stellar mass of the ellipticals without having to undergo a new phase of star formation. The fraction of mass and luminosity contributed by the different stellar populations in the two components further confirms the two-phase assembly mode.
@@
The mass weights from pPXF additionally provide information about the formation ( and quenching ( look-back times, which show that half of the present-day stellar mass in our sample had already assembled Gyr ago, and had undergone quenching around Gyr ago at the latest. The light-weighted properties show similar trends, except for , which indicates a more extended SFH in the outer component, with 90% of the light contribution coming from look-back times as recent as Gyr ago.
@@
In conclusion, our method of decomposing the multiple components of elliptical galaxies has helped trace back the signatures of the physical processes involved in their formation and evolution across cosmic time. This serves as an initial pilot study, which can be extended to a more statistical sample such as the ellipticals in BUDDI-MaNGA (e.g. Jegatheesan et al., 2024) to better constrain the formation pathways. Possible future follow-up work involves capturing the extremely low-surface-brightness stellar halos of elliptical galaxies and modelling their components (Johnston et al., 2018). Additionally, studying a more representative sample of local ellipticals, encompassing a range of masses from dwarf to giant ellipticals, variations in star formation activity, and different environments, would provide a comprehensive understanding of the formation and evolution of these galaxies.